# MathScape: Evaluating MLLMs in multimodal Math Scenarios through a Hierarchical Benchmark

### v3

@@ 1 Introduction
In recent years, there have been advancements in large language models (LLMs) (OpenAI 2023a; Touvron et al. 2023) and MLLMs (Zhao et al. 2023; Wu et al. 2023; Bai et al. 2024). They have shown strong understanding ability among different modalities (Liu et al. 2023b; Bai et al. 2023b).
@@
Among Multimodal Large Language Models (MLLMs), Vision Language Large Models (VLLMs) have demonstrated competitive performance in traditional multimodal tasks, including image classification (Chen et al. 2024), image understanding (Li et al. 2023b, c), and image captioning (Bai et al. 2023b). Furthermore, their advanced language understanding capabilities contribute to strong performance in text-rich tasks, such as visual question answering (Liu et al. 2023b, a) and image-text retrieval (Chen et al. 2024). Recently, VLLMs have also shown significant progress in solving mathematical problems. Therefore, comprehensive benchmarks are essential to evaluate the mathematical abilities of VLLMs. Although several benchmarks, such as MATH-V (Wang et al. 2024a), MathVerse (Zhang et al. 2024), and MathVista (Lu et al. 2023b), have been developed to assess the mathematical capabilities of VLLMs. They primarily focus on a combination of text math problems and image figures. Also, they only use simple metrics and lack effective evaluation for complex or extended responses. Consequently, they face two key challenges:
C1. Insufficient Real-World Data. In previous datasets like MATH-V (Wang et al. 2024a), MathVerse (Zhang et al. 2024), and MathVista (Lu et al. 2023b), the mathematical description was typically provided as text input, while the image contained only figures. This approach doesn’t align well with real-world scenarios, where both the mathematical description and figures are captured together in a single image.
C2. Absence of Effective Evaluation Metrics. In previous datasets (Wang et al. 2024a; Zhang et al. 2024; Lu et al. 2023b), the evaluation was limited to short answers, lacking the ability to assess long-form responses.
@@
Figure 2: An problem example of MathScape. Examples in MathScape are represented by images taken by humans, ensuring a more realistic scenario. Each example will contain a correct answer.
@@ 2.1 Benchmark for Mathematical Evaluation
GSM8K (Cobbe et al. 2021) is a dataset from OpenAI that includes 8.5K high-quality elementary school math word problems, each requiring 2 to 8 steps to solve. These problems primarily involve basic arithmetic operations such as addition, subtraction, multiplication, and division. MATH (Hendrycks et al. 2021) offers a dataset of 12,500 problems sourced from high school math competitions. SuperCLUE-Math (Xu et al. 2024a) is a Chinese benchmark for multi-step reasoning in mathematics, containing over 2,000 problems that require multi-step reasoning and offer natural language solutions. MathBench (Liu et al. 2024b) includes 3,709 math problems ranging from basic arithmetic to college-level questions, covering multiple difficulty levels.
@@
With the rapid advancement of MLLMs, several high-quality benchmarks have emerged to evaluate mathematical problem-solving in visual contexts. MathVista (Lu et al. 2023b) focuses on visual math QA tasks, assessing model performance across various math domains, such as arithmetic and algebra, using visual scenarios. MATH-V (Wang et al. 2024a) is another benchmark that targets multimodal mathematical understanding, with questions primarily sourced from math competitions. MathVerse (Zhang et al. 2024) evaluates MLLMs’ comprehension of visual diagrams using CoT (Chain of Thought) strategies on 2,612 multimodal math problems. CMMU (He et al. 2024) is a large-scale Chinese benchmark for multi-disciplinary, multimodal understanding, featuring questions from college exams and textbooks.
@@ 2.2 MLLMs for Mathematics
The integration of visual knowledge into LLMs has become a pivotal area of research due to the rapid advancements in LLMs. VLLMs combine vision information from vision encoders with LLMs, thus enabling these models to process and interpret visual inputs for various visual tasks  (Liu et al. 2023c; Zhang et al. 2022; Li et al. 2022b) with enhanced accuracy and efficiency. Pioneering frameworks like CLIP  (Radford et al. 2021) leverage contrastive learning on expansive image-caption datasets to align modalities, forming the groundwork for cross-modal comprehension. Various adapters  (Liu et al. 2023b, a; Li et al. 2023b, 2022a; Jian, Gao, and Vosoughi 2023; Lu et al. 2023a) are introduced to further integrate different modalities. For example, LLaVA  (Liu et al. 2023b, a) employs a straightforward MLP to inject the vision information into LLMs. Whereas more complex implementations like the Q-Former in BLIP  (Li et al. 2022a, 2023b) utilize cross-attention to enhance modality integration.
Recent studies  (Wang et al. 2024b; Chen et al. 2023; Liu et al. 2023b, a; Li et al. 2023a) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al. 2023b, a) and ShareGPT4V  (Chen et al. 2023) have shown remarkable advancements in understanding and following complex instructions through instruction tuning.
@@
MAmmoTH (Yue et al. 2023) InternLM-Math (Ying et al. 2024), and ChatGLM-Math (Xu et al. 2024b) are multimodal models specifically tailored for dealing with mathematical questions, incorporating both textual and visual components in their problem design to enhance their ability to handle complex mathematical tasks.
@@ 3.4 Dataset Statistics
Model Algebraic Geometric Equations and Inequality Functions Probability Statistics
@@
4 Experiments and Analysis
4.1 Experimental Setups
Models.
In our evaluation of multimodal LLMs, we focused on both open-source and closed-source models that rank among the top performers on major multimodal LLM leaderboards. This included 11 different types of VLLMs, with a particular emphasis on analyzing the results and performance of the leading models.
  • Closed-source models: GPT4 (OpenAI 2023b), GeminiPro (Reid et al. 2024), Claude-3-Opus, Baichuan-VL (Yang et al. 2023), Qwen-Max (Bai et al. 2023a), Qwn-Plus (Bai et al. 2023a), GLM4V.
  • Open-source models: Deepseek-VL(Lu et al. 2024), LLaVA(Liu et al. 2024a), Yi(Young et al. 2024).
Settings.
@@ 4.1 Experimental Setups
4.2 Performance of Various Models
Question Types
As shown in Table 1, GPT-4V and GPT-4-turbo exhibits the highest accuracy across all question types, with an average of 34.96%, followed by GPT-4-Turbo Vision at 33.92%. While Yi-VL-34B and DeepSeek-V2 achieve good performance among open-source models. We can see the performance of closed-source models achieved better performance than open-source models.
The table shows that models generally perform better on proof questions compared to multiple-choice and solution questions. This suggests that the structured format and clear information in proof questions make them easier for models to handle, while solution questions, which require complex, multi-step reasoning, pose more of a challenge.
Knowledge Points
Educational Stages
Figure 7: The variation of accuracy with answer length.
@@ 5 Challenges and Future Directions
As shown in Section 4, none of the models performed well on our MathScape benchmark. In this section, we discuss the challenges faced by current MLLMs and suggest potential future directions for improving math MLLMs.
@@ 5.2 Future Directions for Math MLLMs
As discussed in section 5.1, it is evident that LLMs have limited mathematical reasoning abilities. To improve the math problem-solving capabilities of MLLMs, it is crucial to develop stronger LLMs with enhanced mathematical reasoning skills.
@@ 6 Conclusion
Appendix A More Case Study
We present additional case studies. As shown in Figure 10 and Figure 11, the model struggles with complex reasoning tasks. This highlights the need to develop stronger reasoning models.
@@ Appendix A More Case Study
Appendix B Prompt for Scoring Answers
@@
Appendix C Visualization of MathScape
C.1 Visualization of Performance of Different Models.
We include additional math samples in MathScape, as illustrated in Figure 13.
Appendix D Reproducibility Checklist
Question1: Includes a conceptual outline and/or pseudocode description of AI methods introduced?
Answer: Yes.
Question2: Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results?
Answer: Yes.
Question3: Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper?
Answer: Yes.
Question4: Does this paper make theoretical contributions?
Answer: Yes.
Question5: All assumptions and restrictions are stated clearly and formally?
Answer: Yes.
Question6: All novel claims are stated formally (e.g., in theorem statements)?
Answer: Yes.
Question7: Proofs of all novel claims are included?
Answer: Yes.
Question8: Proof sketches or intuitions are given for complex and/or novel results?
Answer: Yes.
Question9: Appropriate citations to theoretical tools used are given?
Answer: Yes.
Question10: All theoretical claims are demonstrated empirically to hold?
Answer: Yes.
Question11: All experimental code used to eliminate or disprove claims is included?
Answer: Yes.
Question12: Does this paper rely on one or more datasets?
Answer: Yes.
Question13: A motivation is given for why the experiments are conducted on the selected datasets?
Answer: Yes.
Question14: All novel datasets introduced in this paper are included in a data appendix?
Answer: Yes.
Question15: All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes?
Answer: Yes.
Question16: All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations?
Answer: Yes.
Question17: All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available?
Answer: Yes.
Question18: All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing?
Answer: Yes.
Question19: Does this paper include computational experiments?
Answer: Yes.
Question20: Any code required for pre-processing data is included in the appendix?
Answer: Yes.
Question21: All source code required for conducting and analyzing the experiments is included in a code appendix?
Answer: Yes.
Question22: All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes?
Answer: Yes.
Question23: All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from?
Answer: Yes.
Question24: If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results?
Answer: Yes.
Question25: This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks?
Answer: Yes.
Question26: This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics?
Answer: Yes.
Question27: This paper states the number of algorithm runs used to compute each reported result?
Answer: Yes.
Question28: Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information?
Answer: Yes.
Question29: The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank)?
Answer: Yes.
Question30: This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments?
Answer: Yes.
Question31: This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting?
Answer: Yes.
Figure 13: Math problem samples in MathScape.

### v4

@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across diverse tasks spanning myriad domains OpenAI (2023a); Touvron et al. (2023). Based on LLMs, MLLMs Zhao et al. (2023); Wu et al. (2023); Bai et al. (2024) also show strong understanding ability among different modalities Liu et al. (2023b); Bai et al. (2023b). Among Multimodal Large Language Models (MLLMs), Vision Language Large Models (VLLMs) have demonstrated competitive performance in traditional multimodal tasks, including image classification Chen et al. (2024), image understanding Li et al. (2023b; c), and image captioning Bai et al. (2023b). Furthermore, their advanced language understanding capabilities contribute to strong performance in text-rich tasks, such as visual question answering Liu et al. (2023b; a) and image-text retrieval Chen et al. (2024). Recently, VLLMs have also shown significant progress in solving mathematical problems. Therefore, comprehensive benchmarks are essential to evaluate the mathematical abilities of VLLMs. Although several benchmarks, such as MATH-V (Wang et al., 2024a), MathVerse (Zhang et al., 2024a), and MathVista (Lu et al., 2023b), have been developed to assess the mathematical capabilities of VLLMs. They primarily focus on a combination of text math problems and image figures. Also, they only use simple metrics and lack effective evaluation for complex or extended responses. Consequently, they face two key challenges:
@@
C1. Insufficient Real-World Data. In previous datasets like MATH-V (Wang et al., 2024a), MathVerse (Zhang et al., 2024a), and MathVista (Lu et al., 2023b), the mathematical description was typically provided as text input, while the image contained only figures. This approach doesn’t align well with real-world scenarios, where both the mathematical description and figures are captured together in a single image.
C2. Absence of Effective Evaluation Metrics. In previous datasets Wang et al. (2024a); Zhang et al. (2024a); Lu et al. (2023b), the evaluation was limited to short answers, lacking the ability to assess long-form responses.
@@ 2 Related Work
Figure 2: An example problem from MathScape. Examples in MathScape are represented by images taken by humans, ensuring a more realistic scenario. Each example will contain a correct answer.
@@ 2.1 Benchmark for Mathematical Evaluation
GSM8K Cobbe et al. (2021) is a dataset from OpenAI that includes 8.5K high-quality elementary school math word problems, each requiring 2 to 8 steps to solve. These problems primarily involve basic arithmetic operations such as addition, subtraction, multiplication, and division. MATH Hendrycks et al. (2021) offers a dataset of 12,500 problems sourced from high school math competitions. SuperCLUE-Math Xu et al. (2024) is a Chinese benchmark for multi-step reasoning in mathematics, containing over 2,000 problems that require multi-step reasoning and offer natural language solutions. MathBench Liu et al. (2024b) includes 3,709 math problems ranging from basic arithmetic to college-level questions, covering multiple difficulty levels.
@@
With the rapid advancement of MLLMs, several high-quality benchmarks have emerged to evaluate mathematical problem-solving in visual contexts. MathVista Lu et al. (2023b) focuses on visual math QA tasks, assessing model performance across various math domains, such as arithmetic and algebra, using visual scenarios. MATH-V (Wang et al., 2024a) is another benchmark that targets multimodal mathematical understanding, with questions primarily sourced from math competitions. MathVerse Zhang et al. (2024a) evaluates MLLMs’ comprehension of visual diagrams using CoT (Chain of Thought) strategies on 2,612 multimodal math problems. CMMU He et al. (2024) is a large-scale Chinese benchmark for multi-disciplinary, multimodal understanding, featuring questions from college exams and textbooks.
@@ 2.2 MLLMs for Mathematics
The integration of visual knowledge into LLMs has become a pivotal area of research due to the rapid advancements in LLMs. VLLMs combine vision information from vision encoders with LLMs, thus enabling these models to process and interpret visual inputs for various visual tasks  Liu et al. (2023c); Zhang et al. (2022); Li et al. (2022b) with enhanced accuracy and efficiency. Pioneering frameworks like CLIP  Radford et al. (2021) leverage contrastive learning on expansive image-caption datasets to align modalities, forming the groundwork for cross-modal comprehension. Various adapters  Liu et al. (2023b; a); Li et al. (2023b; 2022a); Jian et al. (2023); Lu et al. (2023a) are introduced to further integrate different modalities. For example, LLaVA  Liu et al. (2023b; a) employs a straightforward MLP to inject the vision information into LLMs. Whereas more complex implementations like the Q-Former in BLIP  Li et al. (2022a; 2023b) utilize cross-attention to enhance modality integration.
Recent studies  Wang et al. (2024b); Chen et al. (2023); Liu et al. (2023b; a); Li et al. (2023a) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  Liu et al. (2023b; a) and ShareGPT4V  Chen et al. (2023) have shown remarkable advancements in understanding and following complex instructions through instruction tuning.
@@
In real-world applications, vision inputs are commonly used to present mathematical problems for models to solve. As a result, it is crucial for Vision-Language Large Models (VLLMs) to demonstrate strong mathematical capabilities. Meidani et al. Meidani et al. (2023) pioneered the use of symbolic data to train a Vision-Language Model (VLM) with mathematical proficiency. Building on this work, UniMath Liang et al. (2023) combined vision, table, and text encoders with LLMs, achieving state-of-the-art (SOTA) performance at the time. Additionally, Huang et al. Huang et al. (2024) succeeded in solving algebraic problems that involved geometric diagrams.
Another noteworthy line of research involves using LLMs to tackle geometric problems. G-LLaVA Gao et al. (2023) fine-tuned LLaVA Liu et al. (2023b) with geometric data, reaching SOTA performance in geometry. Subsequently, MAViS Zhang et al. (2024b) and EAGLE Li et al. (2024) achieved SOTA results by introducing math-specific encoders and amassing large amounts of mathematical data.
@@ 3.4 Dataset Statistics
4 Experiments and Analysis
4.1 Experimental Setups
Models.
In our evaluation of multimodal LLMs, we focused on both open-source and closed-source models that rank among the top performers on major multimodal LLM leaderboards. This included 11 different types of VLLMs, with a particular emphasis on analyzing the results and performance of the leading models. For Closed-source models, we evaluate GPT4 OpenAI (2023b), GeminiPro Reid et al. (2024), Claude-3-Opus, Baichuan-VL Yang et al. (2023), Qwen-Max Bai et al. (2023a), Qwn-Plus Bai et al. (2023a), GLM4V. For Open-source models, we evaluate Deepseek-VLLu et al. (2024), LLaVALiu et al. (2024a), YiYoung et al. (2024).
Settings.
4.2 Performance of Various Models
@@
Question Types
As shown in Table 1, GPT-4V and GPT-4-turbo exhibits the highest accuracy across all question types, with an average of 34.96%, followed by GPT-4-Turbo Vision at 33.92%. While Yi-VL-34B and DeepSeek-V2 achieve good performance among open-source models. We can see the performance of closed-source models achieved better performance than open-source models. The table shows that models generally perform better on proof questions compared to multiple-choice and solution questions. This suggests that the structured format and clear information in proof questions make them easier for models to handle, while solution questions, which require complex, multi-step reasoning, pose more of a challenge.
Knowledge Points
Educational Stages
@@
Model Algebraic Geometric Equations Functions Probability Statistics
@@ 4.4 Answer Length and Accuracy
Figure 5: The variation of accuracy with answer length.
@@ 5 Challenges and Future Directions
As highlighted in Section 4, none of the models achieved strong performance on the MathScape benchmark. In this section, we present several case studies to illustrate the challenges faced by current MLLMs and propose potential future directions for enhancing their mathematical capabilities.
@@ 5.2 Future Directions for Math MLLMs
As outlined in Section 5.1, it is clear that LLMs exhibit limitations in mathematical reasoning. Moreover, all visual information must be processed by the LLM, further constraining its problem-solving capabilities. To enhance the mathematical reasoning proficiency of MLLMs, it is essential to develop more advanced LLMs with stronger mathematical reasoning capabilities.
@@ 6 Conclusion
A. Additional Case Studies
In our case study analysis, we presented three examples that highlight the challenges VLLMs face in solving MathScape problems. These findings underscore the need for more advanced VLLMs capable of tackling mathematical tasks effectively.
B. Visualization of MathScape
In this section, we provide visualizations of the MathScape dataset to offer a clearer and more direct understanding of the benchmark’s structure and content.
C. Prompts for Inference, Answer Extraction, and Scoring
We have summarized the prompts to more clearly demonstrate the process of dataset inference and evaluation, ensuring better reproducibility of the results.
Appendix A Case Study
We present case studies of model responses, highlighting challenges and future directions. As illustrated in Figures 8, 9, and 10, the model faces difficulties with complex reasoning tasks, underscoring the need for more robust reasoning models.
@@ Appendix A Case Study
Appendix B Visualization of MathScape
We include additional math samples in MathScape, translated into English, as shown in Figure 11. Furthermore, we provide examples of human-captured photos within the MathScape dataset.
Figure 11: Math problem samples in MathScape.
Figure 12: Math problem real photo samples in MathScape.
Appendix C Prompt for inference, extracting and Scoring Answers